Each of two parents has the genotype blond/ brown, which consists of the pair of alleles that determine hair color, and each parent contributes one of those alleles to a child. Assume that if the child has at least one blond allele, that color will dominate and the childs hair color will be blond. a. List the different possible outcomes. Assume that these outcomes are equally likely. b. What is the probability that a child of these parents will have the brown/ brown genotype? c. What is the probability that the child will have blond hair color? a. List the possible outcomes. OA. blond/brown and brown/blond yB. blond/blond, blond/brown, brown/blond, and brown/brown OC. blond/blond and brown/brown D. blond/blond, blond/ brown, and brown/brown b. The probability that a child of these parents will have the brown/brown genotype i:s (Round to two decimal places as needed.)
Answer #1

a]

List of possible outcomes are:

blond/blond , blond/brown , brown/blond and brown/brown

b]

The probability that a child of these parents will have the browm/brown genotype is = ratio of favourable outcomes to the total number of outcomes = 1 / 4 = 0.25

c]

Probabilty that the child will have blond hair color = 3 / 4 = 0.75